HIGH RESOLUTION INFRARED STUDY OF THE ν_{4} AND 2ν_{9} BANDS OF ^{11}BF_{2}OH AND ^{10}BF_{2}OH: EVIDENCE OF LARGE AMPLITUDE EFFECTS FOR THE OH-TORSION-BENDING MODES IN THE 9^{2} AND 4^{1} STATES

\underline{A. PERRIN}, E. BERTSEVA, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France; H. BÜRGER, Anorganische Chemie, FBC, Universität, D-42097 Wuppertal, Germany; J. DEMAISON, F. WILLAERT, Laboratoire PhLAM, CNRS, Université de Lille I, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France; A. MASIELLO, T.A. BLAKE, Pacific Northwest National Laboratory, 902 Battelle Boulevard, P.O. Box 999, Richland, WA 99352, USA; A. MARENICH, J.E. BOGGS, Department of Chemistry, University of Texas, Austin, Texas, 78712-1167 USA.

High resolution (2-3 x 10^{−3} cm^{−1}) Fourier transform infrared spectra of gas phase ^{11}B and ^{10}B enriched and natural samples of BF_{2}OH (difluoroboric acid) were recorded at Wuppertal and Richland. Starting from the results of previous studies^{a}, it has been possible to perform the first rovibrational analysis of the ν_{4} (BOH bending) and 2ν_{9} (first overtone of ν_{9}, the OH torsional mode) bands located at 961.5 and 1042.9 cm^{−1}) and at 961.7 and 1043.9 cm^{−1}) for the ^{11}BF_{2}OH and 10BF_{2}OH isotopic species respectively. In addition to various ”classical” vibration-rotation resonances, large amplitude effects were observed for both the 2ν_{9} and ν_{4} bands. These result in a doubling of levels of about 0.004 and 0.003 cm^{−1}) for the 9^{2} and 4^{1} states respectively. The energy level calculations account for these large amplitude torsion-bending perturbations using the IAM- type method^{b}, and for the ”classical” vibration -rotation resonances.

aA.Perrin, M.Carvajal-Zaera, Z.Dutkiewicz, J.M.Flaud, D.Collet, H.Bürger, J.Demaison, F.Willaert, H.Mader, and N.W.Larsen, Mol. Phys. 102 , 1641 (2004), J.Breidung, J.Demaison, J.F.D’Eu, J.M.Flaud, L.Margulès, D.Collet, E.B.Mkadmi, A.Perrin, and W.Thiel, J. Mol. Spectrosc. 228, 7 (2004).

bL.H.Coudert and J.T.Hougen, J. Mol. Spectrosc. 130, 86 (1988)